# Extragradient methods with complexity guarantees for hierarchical variational inequalities

### v1

@@ Appendix B Omitted proofs
Hence,
$\Theta(x|{\mathsf{F}},g,\mathcal{C})=\sup_{y\in\mathcal{C}}\{\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y)\}\leq 0.$
This, together with already obtained inequality $\Theta(x|{\mathsf{F}},g,\mathcal{C})\geq 0$ implies $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$.
Now assume $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$ for $x\in\mathcal{C}$. Then,
$\langle{\mathsf{F}}(x^{\prime}),x-x^{\prime}\rangle+g(x)-g(x^{\prime})\leq 0\hskip 17.00024pt\forall x^{\prime}\in\mathcal{C}.$
Equivalently,
$\langle{\mathsf{F}}(x^{\prime}),x^{\prime}-x\rangle+g(x^{\prime})-g(x)\geq 0\hskip 17.00024pt\forall x^{\prime}\in\mathcal{C}.$
This means $x\in\mathcal{C}$ is a Minty (weak) solution to the hemi-variational inequality with the data $({\mathsf{F}},g)$ over the set $\mathcal{C}$. Let $w\in\mathcal{C}$ be arbitrary and consider v=tw+(1t)xv=tw+(1-t)x for $t\in[0,1]$. Since $\mathcal{C}$ is convex, we have $v\in\mathcal{C}$. It follows
@@
$0\leq\langle{\mathsf{F}}(x),w-x\rangle+g(w)-g(x).$ (B.1)
@@
We now show that if $\mathcal{C}\subset\operatorname{dom}(g)$, and there exists $\epsilon>0$ such that $\mathcal{U}\triangleq B(x,\epsilon)\cap\mathcal{C}=B(x,\epsilon)\cap\operatorname{dom}(g)$, then (B.1) holds also for $w\in\operatorname{dom}(g)$. To that end, assume to the contrary, there exits $z\in\operatorname{dom}(g)$ for which
$\langle{\mathsf{F}}(x),z-x\rangle+g(z)-g(x)<0.$
Then, by convexity of $\operatorname{dom}(g)$ and definition of $\mathcal{U}$ there exists $\lambda>0$ small enough for which $w=x+\lambda(z-x)\in\mathcal{U}$ and by convexity
@@
thus arriving at a contradiction. Hence, $\langle F(x),z-x\rangle+g(z)-g(x)\geq 0$ for any $z\in\operatorname{dom}(g)$ and moreover for any $z\in\mathcal{Z}$, which finishes the proof that $x$ is a solution to (2.3). $\blacksquare$
@@
Pick $z\in\operatorname{dom}(g)$. Without loss of generality, assume $z\notin\mathcal{S}$. Let $\bar{z}=\Pi_{\mathcal{S}}(z)$, so that there exists $p\in\operatorname{\mathsf{NC}}_{\mathcal{S}}(\bar{z})$ satisfying $p=z-\bar{z}$. Hence, $p^{\ast}\triangleq\frac{z-\bar{z}}{\norm{z-\bar{z}}}$ is a unit norm element of $\operatorname{\mathsf{NC}}_{\mathcal{S}}(\bar{z})$ satisfying
$\langle p^{\ast},z-\bar{z}\rangle=\norm{z-\bar{z}}.$
@@
$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+g_{1}(z)-g_{1}(z^{*})\geq\langle-p^{*},z-z^{*}\rangle\hskip 17.00024pt\forall z\in\mathcal{Z}.$ (B.2)
Take $z\in\mathcal{Z}$ and let $\hat{z}=\Pi_{\mathcal{S}_{2}}(z)$ be the orthogonal projector of $z$ onto $\mathcal{S}_{2}$. Thus, $\langle p^{*},\hat{z}-z^{*}\rangle\leq 0$, resulting in
$\langle p^{*},\hat{z}-z\rangle\leq\langle p^{*},z^{*}-z\rangle,$
which implies when combined with (B.2)
$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+$ $g_{1}(z)-g_{1}(z^{*})\geq\langle p^{*},\hat{z}-z\rangle$
$\geq-\norm{p^{*}}\cdot\norm{\hat{z}-z}=-\norm{p^{*}}\operatorname{dist}(z,\mathcal{S}_{2}).$
Hence, for all compact $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})$ with $z^{*}\in\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$, we conclude $\Theta_{\rm Opt}(z|\mathcal{U}_{1}\cap\mathcal{S}_{1})\geq-B_{\mathcal{U}_{1}}\operatorname{dist}(z,\mathcal{S}_{2})$, where $B_{\mathcal{U}_{1}}=\sup_{\mathcal{U}_{1}\cap\mathcal{S}_{1}}\norm{p^{*}}$.

### v2

@@ Appendix C Omitted proofs
The convexity of $x\mapsto\Theta(x|{\mathsf{F}},g,\mathcal{C})$ follows from the fact that it is a supremum of convex functions. Since $\mathcal{C}$ is compact, this function is well-defined.
@@
Hence, $\Theta(x|{\mathsf{F}},g,\mathcal{C})=\sup_{y\in\mathcal{C}}\{\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y)\}\leq 0$. This, together with already proved inequality $\Theta(x|{\mathsf{F}},g,\mathcal{C})\geq 0$ implies $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$.
Now assume $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$ for $x\in\mathcal{C}$. Then, $\langle{\mathsf{F}}(x^{\prime}),x-x^{\prime}\rangle+g(x)-g(x^{\prime})\leq 0$ for all $x^{\prime}\in\mathcal{C}.$ Equivalently, $\langle{\mathsf{F}}(x^{\prime}),x^{\prime}-x\rangle+g(x^{\prime})-g(x)\geq 0$ for all $x^{\prime}\in\mathcal{C}.$ This means $x\in\mathcal{C}$ is a Minty (weak) solution to the $\operatorname{HVI}({\mathsf{F}},g)$ over the set $\mathcal{C}$. Let $w\in\mathcal{C}$ be arbitrary and consider v=tw+(1t)xv=tw+(1-t)x for $t\in[0,1]$. Since $\mathcal{C}$ is convex, we have $v\in\mathcal{C}$. It follows
@@
$0\leq\langle{\mathsf{F}}(x),w-x\rangle+g(w)-g(x).$ (C.1)
@@
[…]
@@
thus arriving at a contradiction. Hence, $\langle F(x),z-x\rangle+g(z)-g(x)\geq 0$ for any $z\in\operatorname{dom}(g)$ and moreover for any $z\in\mathcal{Z}$, which finishes the proof that $x$ solves (2.3). $\blacksquare$
@@
$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+g_{1}(z)-g_{1}(z^{*})\geq\langle-p^{*},z-z^{*}\rangle\hskip 17.00024pt\forall z\in\mathcal{Z}.$ (C.2)
Take $z\in\mathcal{Z}$ and let $\hat{z}=\Pi_{\mathcal{S}_{2}}(z)$. Thus, $\langle p^{*},\hat{z}-z^{*}\rangle\leq 0$, resulting in $\langle p^{*},\hat{z}-z\rangle\leq\langle p^{*},z^{*}-z\rangle,$ which implies when combined with (C.2)
$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+$ $g_{1}(z)-g_{1}(z^{*})\geq\langle p^{*},\hat{z}-z\rangle\geq-\norm{p^{*}}\cdot\norm{\hat{z}-z}=-\norm{p^{*}}\operatorname{dist}(z,\mathcal{S}_{2}).$
Hence, for all compact $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})$ with $z^{*}\in\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$ and corresponding $p^{*}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{*})$, we conclude $\Theta_{\rm Opt}(z|\mathcal{U}_{1}\cap\mathcal{S}_{1})\geq-B_{\mathcal{U}_{1}}\operatorname{dist}(z,\mathcal{S}_{2})$, where $B_{\mathcal{U}_{1}}\triangleq\norm{p^{*}}$.